OpenStudy (anonymous):

Would anyone like to be my study buddy? Please I really need one. PLEASE I WILL FAN. Message me if you would like to be my study buddy.

OpenStudy (anonymous):

ill try!

OpenStudy (gabebae):

math?

OpenStudy (anonymous):

Everything!

OpenStudy (gabebae):

What do you need help with now?

OpenStudy (anonymous):

Right now I am studying for my mid-unit quiz. That quiz is about adding and subtracting fracions with unlike and like denominators.

OpenStudy (gabebae):

Do you need help with questions or the entire lesson?

OpenStudy (anonymous):

It is math.

OpenStudy (gabebae):

I know

OpenStudy (anonymous):

I just need to sort of run through it.

OpenStudy (anonymous):

Hi lady ginger you never replied to my message.

OpenStudy (anonymous):

Why?

OpenStudy (gabebae):

When you're subtracting and adding fractions with like denominators you just add/subtract straight through Example: \[\frac{ 4 }{ 9 } + \frac{ 3 }{ 9 } = \frac{ 7 }{ 9 }\] When you're subtracting and adding fractions with different denominators you first have to find the least common factor so you could get the same denominator Example: \[\frac{ 1 }{ 2 } + \frac{ 3 }{ 10 }\] the common denominator would be 10

OpenStudy (anonymous):

Ok thanks.

OpenStudy (gabebae):

Hope it helps?

OpenStudy (gabebae):

Khan Academy can you help you more

OpenStudy (anonymous):

I know Khan Academy I go there all the time. Thanks it helped.